# WarmSwap: Sharing Dependencies for Accelerating Cold Starts in Serverless Functions

### v1

@@ 2.3 Performance Analysis of the Cold Start for Current AWS Lambda Functions
Figure 3: Cold-start latency breakdown for serverless functions in a traditional AWS Lambda function (not WarmSwap).
@@
Figure 3 shows the breakdown of intermediate times for cold starts in this analysis. The entire cold-start time can be broken down into to following parts:
@@
From Figure 3, the dependency initialization dominates the cold start of AWS Lambda functions. In “lr_serving”, “cnn_serving”, and “rnn_serving”, the dependency initialization occupies $43\%$, $74\%$ and $68\%$ of the cold-start overhead, respectively. The time for the dependency initialization component is even higher than that for the execution component for the three AWS Lambda functions. The dependency initialization component is a major problem for serverless functions in production. This problem occurs frequently because the packages they need are so common.
Note that the time for the container component is almost constant (about $0.5~{}s$), even for containers that have different image sizes. This shows that AWS has already optimized the time for the container component, by improving the AWS caching and file systems [5, 1].
In summary, the time of the dependency initialization component often dominates the time for the other components. Therefore, improving the dependency initialization component overhead is important both to serverless providers and to users of the current production cloud. This motivates a novel method focusing on the dependency initialization overhead, as seen in Section 3.
@@
Figure 4: System architecture and workflow: (a) Architectureal overview of WarmSwap and its dependency pool. (b) In the setup phase, the user registers a serverless function, and WarmSwap generates the corresponding image in the dependency pool. (c) In the runtime phase, pages are copied from the image between provider worker nodes and user function containers.
3 WarmSwap Design
This section introduces WarmSwap, a prototype system for enabling “live dependency image” on serverless platforms. WarmSwap is motivated by the goal of improving the cold start by skipping the dependency loading phase. WarmSwap enables a serverless function container to seamlessly migrate the dependency image of its pre-initialized software dependencies from a provider’s worker node. For example, assume there is a serverless function based on Python and numpy. The function can use pre-initialized memory pages of Python and numpy through migration. This eliminates the otherwise large runtime overhead of first initializing Python and numpy.
WarmSwap leverages the power of a shared dependency pool, enabling significant performance gains for providers. Even though individual functions differ, multiple users whose functions have a common dependency can benefit from the same pre-initialized dependency image from the pool. The cloud provider can then support such a dependency pool for all serverless users. In general, the pool contains several pre-initialized dependency images for popular runtimes and packages (e.g., Python + numpy, Python + torch, etc.).
We first provide an overview of the system’s architecture in Section 3.1, highlighting the key components and their interactions. Next, we detail the implementation of each WarmSwap component, including the generation of the pool of dependency images in Section 3.2 and the design choices of migration in Section 3.3. Finally, Section 3.4 describes the implementation.
3.1 Design Overview
WarmSwap is designed to reduce dependency loading overhead for serverless functions. The key idea is to migrate pre-initialized dependencies instead of starting them from scratch. Here, migrate means to copy a process from the host (which could be a local or remote node) to the user container over the network. The system design is based on the characteristics of such dependencies. A container for a serverless function usually contains both user-specific code and a middleware platform with its own package dependencies. We refer to this middleware as a serverless software dependency. These containers generally share three characteristics, described below, which determine the design and performance of the WarmSwap workflow.
@@ 3.1 Design Overview
Although there are numerous serverless software dependencies, and applications can use different combinations of dependencies, many applications are based on a small number of popular packages (e.g., NumPy, sklearn, torch). Providing migration services for those popular packages can benefit complex serverless functions with large dependencies.
@@
Most applications do not need all functions from a package, which means they only need access to small parts of the dependency memory pages. In this situation, it is unnecessary to load all the memory pages before the execution of user functions.
@@
Keeping runtime and software dependencies in the memory is much more lightweight than keeping the whole container. Thus, the provider can afford to maintain a pool of dependency images on their work node.
@@
Figure 4a describes the components of WarmSwap on the provider work node and function container. On the provider worker node, a Dependency Manager orchestrates the management of in-memory or "live" dependency images, which are images stored in memory that represent a checkpoint of pre-initialized software dependencies (e.g., Python + numpy). It is the central hub for communication with user function containers, facilitating migration services. The Dependency Manager maintains metadata about existing live dependency images and manages the memory pages associated with them. When a migration request is initiated by a serverless container, the Dependency Manager accepts the request, transfers the essential metadata required for process restoration, and provides page transfer services to the serverless function container.
@@
In the container for the user function, a Migration Client orchestrates the migration of memory pages on the user side. Upon startup of the container, it initiates a migration request for the Dependency Manager. Upon receiving confirmation, it starts the Dependency Loader to orchestrate the restoration of the runtime and the loading of dependency images, guided by the provided metadata.
@@
Rather than loading all memory pages immediately, the Dependency Loader restores the memory pages of the dependency by lazy migration. The Dependency Loader restores the dependency without loading the actual memory pages; instead, it initiates a Page Server to handle this task later. When the user functions execute and access memory pages of the dependency, it triggers page faults, and the Page Server handles them. This Page Server communicates with the dependency pool to efficiently load specific pages whenever required by the process.
3.2 Generation of Dependency Images
@@
Figure 4b shows the setup phase of WarmSwap, which dumps a serverless software dependency into a dependency image in the dependency pool. The serverless software dependency process (refer to the target process in the following paragraphs) in this example is a Python script importing numpy and then executing the user handler. We initiate the dumping of the process after numpy has been loaded but before the import and execution of the user code (the user handler in Figure 4a). This dump timing ensures that the migrated process can support different user handlers and execute them immediately after migration.
@@
The target process is divided into two distinct components during the dumping process: process metadata, preserved within Dependency Metadata; and hot memory pages, residing securely within the Dependency pool. The process metadata encompasses essential information for restoring the process framework, including file descriptors, memory segment sizes, and other relevant configuration details. The dependency pool preserves the full memory pages of the target process within its memory domain. To enable efficient page transfer, the Dependency Manager creates a dedicated page server process on the source for each target process. The source page server uses the network to communicate with page servers on the destination that need the target process.
To further accelerate the process generation and optimize resource utilization, we employ checkpoint images on disk. When dumping target processes, we create checkpoint images in order on disk to further optimize the generation process. With images on disk available, we do not need to execute the Python script from the beginning. Instead, we can directly restore the checkpoint and then create live dependency images. This is useful for those large and less popular target processes because the provider can save costs by putting them on disk instead of memory.
@@
WarmSwap requires only a modest pool of dependency images, because of the commonality and lightweight characteristics of serverless software dependencies. Thus, cloud providers only undertake limited costs associated with dumping and maintaining dependent images.
3.3 Migration of Dependency Images
Figure 4c describes the runtime details of initializing a WarmSwap container. Due to the sparsity of the memory pages required by a dependency image, we leverage a strategy known as lazy migration. This approach dynamically transfers pages only as needed, significantly reducing overhead and enhancing performance.
@@ 3.3 Migration of Dependency Images
The first step of migration is to transport process metadata to the destination (i.e., to the serverless function container). After receiving the migration request from the migration client, the dependency manager looks up the dependency metadata for the requested dependency image. If the dependency image exists, the dependency manager transports process metadata to the dependency loader.
The second step of migration is to start a page server, which handles the “Page Faults” of the dependency to be loaded by using the “userfaultfd” feature of Linux. The page server accesses the dependency pool through the network, to load those pages that actually are needed by the serverless function.
@@
However, the dependency loader does not load all of the memory pages in the dependency image. Instead, it starts a page server that connects to a specific dependency image inside the dependency pool on the worker node. Thus, when the process accesses specific memory pages, it triggers a “Page Fault”, which the page server handles using “userfaultfd”.
For handling the “Page Fault”, WarmSwap could implement either of two possible design choices:
@@
Bulk restore: When handling the first “Page Fault”, the Page Server first transfers the required pages. And then, in the background, it loads the remaining pages in parallel until another page fault occurs. In this case, the Page Server needs to access the dependency pool only once. Therefore, the execution will not be impacted by further page faults and remote communication. Although transferring all pages increases communication overhead, this should not impact the user functions because the Page Server can work in parallel with the user functions.
WarmSwap chooses bulk restore because it generally benefits all functions without impacting the warm starts (see Section 4.4). In contrast, lazy restore only works well in fewer page fault cases. In a third possibility, one might record the subset of pages that the serverless functions need and only transfer them. For example, previous work [25] records and pre-fetch hot pages for each target serverless function. However, recall that this method involves function-specific tuning, which does not fulfill the requirements of WarmSwap.
3.4 Implementation
We implement a prototype of WarmSwap with AWS Lambda containers and CRIU [7] (Checkpoint and Restore In Userspace). WarmSwap uses AWS Lambda container templates (Version Python 3.9) to build serverless functions and collect performance results. The dependency loader is embedded into the boot process of an AWS Lambda container. When a serverless container receives invocation, it first starts the dependency loader instead of the original “bootstrap” script. Then the dependency loader loads dependency images into this container and continually executes the AWS Lambda Runtime Interface [4] and the user code.
Note that WarmSwap is a provider-side implementation. It was necessary to simulate WarmSwap on an AWS EC2 node instead of the online AWS Lambda environment. Nevertheless, all experiments employ a standard AWS Lambda container and simply execute the container with the EC2 environment. The EC2 environment was required because AWS Lambda does not allow the user to control the placement of serverless functions on nodes. In particular, it cannot guarantee that a serverless function will be placed on the same node as its associated pool of dependency images.
The prototype modifies CRIU (Version 3.9) to implement live migration of dependency images. We add two new features supporting the design described in Section 3.2 and Section 3.3.
First, we modified the “criu dump” command to support the dependency pool. This ensures that the “criu dump” command will always keep the live dependency image in memory (not killed after migration) and managed by the dependency manager of WarmSwap. Additionally, we extend the live dependency image to accept multiple requests.
Second, we modified the “criu pageserver” command to support the lazy restore policy described in Section 3.3. Note that there is a “--lazy page” parameter of the “criu pageserver” command. However, this parameter migrates all pages upon the execution start. We need to implement WarmSwap’s lazy restore policy: loading pages one-by-one. Similarly, CRIU’s --lazy page parameter does not fully support WarmSwap’s bulk restore policy. Some smaller modifications were also required for this policy.
Finally, it is important that the Python runtime and packages are installed in the same path and with the same version on both the worker node and the container. Otherwise, the “criu restore” cannot find the necessary files and reconnect the file descriptors.
4 Evaluation
We first present the experimental setup in Section 4.1, including the experimental platform and benchmarks for serverless functions. The implementation was described in Section 3.4. The evaluation answers the following research questions:
  • RQ1: How well does WarmSwap improve the cold-start performance of different serverless functions? (Section 4.2)
  • RQ2: How does each component in WarmSwap contribute to the cold start of different serverless functions? (Section 4.3)
  • RQ3: How well does live migration in WarmSwap improve the cold-start execution compared with standard migration using simple checkpoint images? (Section 4.4)
  • RQ4: How do different migration policies (lazy restore and bulk restore) work in WarmSwap? (Section 4.4)
  • RQ5: How well does WarmSwap perform when sharing dependency images across multiple serverless functions? (Section 4.5)
4.1 Experimental Setup
We use a “r5b.large” AWS EC2 machine with two vCPUs, 16 GB memory, and AWS EBS 200 GB SSD, to conduct our experiments. The O/S is Amazon Linux with Linux kernel 5.10.199, and the gcc version is 10.5.0. As described in Section 3.4, an AWS EC2 node is required to simulate a provider-side WarmSwap implementation, but standard AWS Lambda server containers are used in all experiments. Moreover, all serverless containers are built based on the “public.ecr.aws/lambda/python:3.9” template. The serverless containers are started with “--privileged” to allow the execution of CRIU.
All experiments use Python-based functions from the FunctionBench [12, 13] serverless benchmark suite. Recall that the middleware packages needed by these functions are described in Table 1 in Section 2.3. The purposes of these serverless functions are as follows:
  • helloworld: Minimal function with no dependency packages.
  • pyase: Text encryption with an AES block-cipher.
  • chameleon: HTML table rendering.
  • json_dumps_load: Serialization and de-serialization of JSON.
  • lr_serving: Serving a logistic regressions model.
  • cnn_serving: Serving an image classification model based on CNN.
  • rnn_serving: Serving a sequence generation model based on RNN.
Functions needing model parameters (namely, lr _serving, cnn _serving, rnn _serving) download their parameters from AWS S3 when executing cold starts. The dependency images created for each serverless function are based on Python and Python packages described in Table 1.
4.2 Comparing WarmSwap and Baseline
To answer RQ1, we evaluate the cold-start latency of WarmSwap and the Baseline with all functions described in Section 4.1. Baseline refers to a traditional AWS Lambda container running offline. The Baseline runs the same containers that were built for online AWS Lambda in Section 2.3 on an EC2 machine. In this experiment, WarmSwap uses bulk restore (see Section 3.3) because it has the best performance. These serverless containers directly boot the AWS Lambda Runtime Interface during cold starts and use a Python environment and packages pre-installed inside the serverless container.
@@ 4.2 Comparing WarmSwap and Baseline
Figure 5: Comparison of WarmSwap versus Baseline. The software dependencies for each function are specified in Table 1 of Section 2.3. Note that lr_serving, cnn_serving, and rnn_serving all have much larger dependency images. For this reason, WarmSwap is shown to be much faster on cold starts than the Baseline in these cases, by a ratio of 1.2, 1.8, and 2.2. Further, there is almost no degradation of performance during warm starts.

### v3

@@ II-C Performance Analysis of the Cold Start for Current AWS Lambda Functions
Figure 3: Cold-start latency breakdown for serverless functions in a traditional AWS Lambda function (not HotSwap).
This section analyzes the performance of cold-start overhead in a production cloud on AWS [23, 22]. The goal is to analyze which components of the current AWS Lambda architecture still have room for optimization.
For analysis, we build serverless function containers to create AWS Lambda functions, and then invoke them by a script. All Lambda functions are configured with a 10 GB memory limitation to maximize the CPU resources. To collect cold-start times, we delete and recreate lambda functions after each run. There are several timers in the analysis script and the function containers to track the time of each cold-start process. The entire time for the cold start is calculated from the script that sends the request to the serverless function and receives the result.
Note that the caching system of AWS Lambda would cause bias in a naive experimental design. For example, if the user uploads a new function container, the first cold start will be extremely slow as compared to the following ones. Moreover, if the functions are frequently invoked, AWS upgrades the function to a different cache that will improve its cold starts. (See “AWS three-tier caching” in [22].) The improved caching scheme of AWS Lambda continues to be used even when the function is deleted and then recreated. In order to achieve reproducible numbers with lower variance in our own experiments on AWS Lambda, we upload a container and then create and delete a Lambda function many times, until the cold-start results are stable.
TABLE I: Serverless functions adopted from FunctionBench [24].
Serverless Functions Container Size (MB) Major Dependencies Function Purpose
helloworld 180.00 None Minimal function
json_dumps_load 180.00 urllib, json Json Serialization
pyase 180.11 pyaes, string Text encryption
chameleon 180.42 chameleon HTML rendering
lr_serving 379.52 sklearn, pandas Logistic regressions
cnn_serving 1386.73 NumPy, keras Image classification
rnn_serving 5602.46 NumPy, torch Sequence generation
Table I presents serverless functions from a representative serverless benchmark suite called FunctionBench [24, 25]. We have chosen FunctionBench for ease of comparison with previous work [17]. All these functions are written in Python and they use some representative Python packages like “NumPy”, “torch”, “keras”, and “sklearn”. Table I lists both packages and the size of specific function container images. Installing large packages also significantly increases the image size of the function containers.
@@
Figure 3 shows the breakdown of intermediate times for cold starts in this analysis. The entire cold-start time can be broken down into the following parts:
@@
From Figure 3, the dependency initialization dominates the cold start of AWS Lambda functions. In “lr_serving”, “cnn_serving”, and “rnn_serving”, the dependency initialization occupies $43\%$, $74\%$, and $68\%$ of the cold-start overhead, respectively. Dependency initialization takes more time than execution for all three AWS Lambda functions. It remains a major bottleneck for serverless functions in production, especially since many rely on widely used packages.
Note that the time for container creation is almost constant (about $0.5~{}s$), even for containers that have different image sizes. This shows that AWS has already optimized the time for the container creating, by improving the AWS caching and file systems [22, 23].
In summary, the time for dependency initialization often dominates the cold-start time in AWS Lambda. Therefore, improving the dependency initialization overhead is important both to serverless providers and to users of the current production cloud. This motivates a novel method focusing on the dependency initialization overhead, as seen in Section III.
@@
Figure 4: System architecture and workflow: (a) Architectural overview of HotSwap and its dependency pool. (b) In the setup phase, the user registers a serverless function, and HotSwap generates the corresponding image in the dependency pool. (c) In the runtime phase, pages are copied from the image between provider worker nodes and user function containers.
III HotSwap Design
This section introduces HotSwap, a prototype system for cross-function cold-start optimization. HotSwap leverages the power of a shared dependency pool, enabling significant performance gains for providers. A single global shared dependency pool for all users is implemented for cross-function optimization. Hence, a single pre-initialized dependency image from the pool can benefit multiple users whose functions share a single common dependency. In general, the pool contains several pre-initialized dependency images for popular runtimes and packages (e.g., NumPy, torch, etc.). Thus, the cloud provider supports such a dependency pool for all users.
HotSwap enables a serverless function container to seamlessly migrate the dependency image of its pre-initialized software dependencies from a provider’s worker node. For example, assume there is a serverless function based on NumPy. The function can use pre-initialized memory pages of NumPy through migration. This eliminates the otherwise large runtime overhead of first initializing NumPy.
III-A Motivation for the Design of HotSwap
HotSwap gains its performance advantages by exploiting three characteristics of the workloads seen by public cloud providers in their distribution of containerized serverless functions. Those three characteristics are:
@@ III-A Motivation for the Design of HotSwap
Although there are numerous serverless software dependencies, and applications can use different combinations of dependencies, many applications are based on a small number of popular packages [19] (e.g., NumPy, torch).
@@
Most applications do not use all functions from a package during execution [17]. Thus, copying all memory pages to the function container beforehand is unnecessary.
@@
Keeping runtime and software dependencies in the memory is much more lightweight than keeping the whole container like [19]. Thus, the provider can afford to maintain a pool of dependency images.
III-B Design Overview
HotSwap is designed to reduce dependency loading overhead for serverless functions. The key idea is to migrate pre-initialized dependencies instead of starting them from scratch. Here, migrate means to copy a process from the host (which could be a local or remote node) to the user container over the network. The system design is based on the characteristics of such dependencies. A container for a serverless function usually contains both user-specific code and its software dependencies.
@@ III-B Design Overview
Figure 4a describes the components of HotSwap on the provider worker node and function container. On the provider worker node, a Dependency Manager orchestrates the management of in-memory dependency images, which are images stored in memory that represent a checkpoint of pre-initialized software dependencies. The Dependency Manager maintains metadata about existing live dependency images and manages the memory pages associated with them. It is the central hub for communication with user function containers, facilitating migration services. When a migration request is initiated by a serverless container, the Dependency Manager accepts the request, transfers the essential metadata required for process restoration, and provides page transfer services to the serverless function container.
@@
In the container for the user function, a Migration Client orchestrates the migration of memory pages on the user side. Upon startup of the container, it initiates a migration request to the Dependency Manager. Upon receiving confirmation, it starts the Dependency Loader to orchestrate the restoration of the runtime and the loading of dependency images, guided by the provided metadata.
@@
Rather than loading all memory pages immediately, the Dependency Loader restores the memory pages of the dependency by lazy migration. In lazy migration, instead of the actual memory pages, the Dependency Loader initializes a Page Server to handle this task later. When the user functions execute and access memory pages of the dependency, it triggers page faults, and the Page Server handles them. This Page Server communicates with the dependency pool to efficiently load memory pages.
III-C Generation of Dependency Images
@@
Figure 4b shows the setup phase of HotSwap, which dumps a serverless software dependency into a dependency image in the dependency pool. The serverless software dependency process in this example is a Python script importing NumPy and then executing the user handler. We initiate the dumping of the process after NumPy has been loaded but before the import and execution of the user code (the user handler in Figure 4a). In general, other shared dependency images might also be used besides NumPy. The corresponding user handler will then execute immediately after that image has been migrated.
@@
The dependency process is divided into two distinct components during the dumping process: (i) process metadata, preserved within Dependency Metadata; and (ii) hot memory pages, residing securely within the Dependency Pool. The process metadata encompasses essential information for restoring the process framework, including file descriptors, memory segment sizes, and other relevant configuration details. The dependency pool preserves the full memory pages of the dependency process within its memory domain. The process metadata, combined with its memory pages, is referred to as a dependency image in this work. To enable efficient page transfer, the Dependency Manager creates a source-side page server for each dependency image, which communicates over the network with destination page servers.
@@
HotSwap requires only a modest pool of dependency images, because of the commonality and lightweight characteristics of serverless software dependencies. Thus, cloud providers only undertake limited costs associated with dumping and maintaining dependency images.
III-D Migration of Dependency Images
@@
The first step of migration is to transport the process metadata to the destination (i.e., to the serverless function container). After receiving the migration request from the migration client, the dependency manager looks up the dependency metadata for the requested dependency image. If the dependency image exists, the dependency manager transports process metadata to the dependency loader.
The second step of migration is to start a page server, which handles the “page faults” of the dependency to be loaded by using the “userfaultfd” feature of Linux. The page server accesses the dependency pool through the network, to load those pages that are actually needed by the serverless function.
@@
For handling the “page fault”, HotSwap could implement either of two possible policies:
@@
Bulk restore: When handling the first “page fault”, the Page Server first transfers the required pages. And then, in the background, it loads the remaining pages in parallel. In this case, the Page Server needs to access the dependency pool only once. Therefore, the execution will not be impacted by further page faults and remote communication. Although transferring all pages increases communication overhead, this should not impact the user functions because the Page Server can work in parallel with the user functions.
HotSwap chooses bulk restore because it generally benefits all functions without impacting the warm starts (see Section IV-D). In contrast, lazy restore works well only in more specialized page fault cases.
@@
Figure 5: Comparison of HotSwap versus Baseline. The software dependencies for each function are specified in Table I of Section II-C. Note that lr_serving, cnn_serving, and rnn_serving all have much larger dependency images. For this reason, HotSwap is shown to be much faster on cold starts than the Baseline in these cases, by a ratio of 1.2, 1.8, and 2.2. Further, there is almost no degradation of performance during warm starts.
III-E Implementation
We implement a prototype of HotSwap with AWS Lambda containers and CRIU (Checkpoint and Restore In Userspace) [27]. HotSwap uses AWS Lambda container templates (Python 3.9) to build serverless functions and collect performance results. The dependency loader is embedded into the boot process of an AWS Lambda container. When a serverless container receives an invocation, it first starts the dependency loader instead of the original “bootstrap” script. Then the dependency loader loads the dependency image into the container and continually executes the AWS Lambda Runtime Interface [26] and the user code.
HotSwap is simulated on an AWS EC2 node instead of the online AWS Lambda environment. This was necessary, since HotSwap is a provider-side implementation, and it is not possible to modify AWS Lambda. Nevertheless, all experiments employ a standard AWS Lambda container and simply execute the container with the EC2 environment. The EC2 environment was required because AWS Lambda does not allow the user to control the placement of serverless functions on nodes.
The prototype modifies CRIU (Version 3.9) to implement live migration of dependency images. We add two new features supporting the design in Section III-C and Section III-D.
First, we modified the “criu dump” command to support the dependency pool. This ensures that the “criu dump” command will always keep the live dependency image in memory (instead of being killed after migration) and managed by the dependency manager of HotSwap.
Second, we modified the “criu pageserver” command to support the lazy restore policy described in Section III-D. We extend the CRIU pageserver to accept multiple requests. Note also that there is a “--lazy page” parameter of the “criu pageserver” command. However, this parameter migrates all pages upon the execution start. We need to implement HotSwap’s lazy restore policy: loading pages one by one. Similarly, CRIU’s --lazy page parameter does not fully support HotSwap’s bulk restore policy. Some smaller modifications were also required for this policy.
We will release our configuration files and prototype code online as open source, to aid in reproducibility.
IV Evaluation
We first present the experimental setup in Section IV-A, including the experimental platform and benchmarks for serverless functions. The evaluation answers the following research questions:
  • RQ1: How well does HotSwap improve the cold start of different serverless functions? (Section IV-B)
  • RQ2: How does each HotSwap component affect cold starts across serverless functions? (Section IV-C)
  • RQ3: How well does live migration in HotSwap improve the cold-start execution compared with standard migration using simple checkpoint images? (Section IV-D)
  • RQ4: How do different migration policies (lazy restore and bulk restore) work in HotSwap? (Section IV-D)
  • RQ5: How does HotSwap compare to PreBaking [13, 14] when sharing dependency images across multiple serverless functions in a real-world trace? (Section IV-E)
IV-A Experimental Setup
We use “r5b.large” AWS EC2 machines with two vCPUs, 16 GB memory, and AWS EBS 200 GB SSD, to conduct our experiments. The OS is Amazon Linux with Linux kernel 5.10.199, and the gcc version is 10.5.0. As described in Section III-E, an AWS EC2 node is required to simulate a provider-side HotSwap implementation, but standard AWS Lambda server containers are used in all experiments. Moreover, all serverless containers are built based on the “public.ecr.aws/lambda/python:3.9” template. The serverless containers are started with “--privileged” to allow the execution of CRIU.
All experiments use Python-based functions from the FunctionBench [24, 25] serverless benchmark suite. Recall that the middleware packages needed by these functions and other details are described in Table I in Section II-C.
Functions needing model parameters (namely, lr _serving, cnn _serving, rnn _serving) download their parameters from AWS S3 when executing cold starts. The dependency images created for each serverless function are based on Python and Python packages described in Table I.
IV-B Comparing HotSwap and Baseline
To answer RQ1, we evaluate the cold-start latency of HotSwap and the Baseline with all functions described in Section IV-A. Baseline refers to a traditional AWS Lambda container running offline. The Baseline runs the same containers that were built for online AWS Lambda in Section II-C on an EC2 machine. In this experiment, HotSwap uses bulk restore (see Section III-D) because it has the best performance. These serverless containers directly boot the AWS Lambda Runtime Interface during cold starts and use a Python environment and packages pre-installed inside the serverless container.